# Gravity-Informed Deep Learning Framework for Predicting Ship Traffic Flow and Invasion Risk of Non-Indigenous Species via Ballast Water Discharge

### v1

@@ Discussion
The ability to assess and predict the risks of NIS invasion through ballast water relies heavily on shipping fluxes. In order to improve BWRA, we have introduced a new approach called the Transformer Gravity model. Our Transformer Gravity model is built based on a stack of multi-head attention blocks. It incorporates features originating from the gravity-informed model, including the shipping fluxes, origin and destination locations, and the geographical distance between them. We enriched the feature set by integrating additional factors, including the international bilateral trade information and the graph metrics analyzed from the global shipping network, to enhance the model’s capability for predicting ship traffic flows. We have validated our approach through a comprehensive comparison with established gravity models and a range of machine-learning regression techniques to highlight its advancements.
Our findings established the superiority of the Transformer Gravity model across several performance indicators. Notably, the model achieved the highest CPC with minimal variance across different data folds, indicating its robustness and adaptability to new datasets. The best average CPC in the cross-validation set was 0.864 for the 3 and 5-layer models. Besides, the 3-layer one demonstrated the lowest mean error, 0.080, and the highest mean correlation, 0.977. Variations in performance among baseline models highlighted their inadequacies in dealing with diverse data scenarios, thus reinforcing our model’s resilience.
Our study has limitations that can be improved with future and further research. For instance, our model tends to be over-optimistic in targeting high-risk ship traffic predictions, and there might be underlying features that can contribute more to forecasting high-risk trips. Seasonality of environmental variables can affect the results of environmental distance calculation, and using environmental data collected from a smaller temporal scale can reflect the evaluated risk more precisely. Our study did not consider ecological barriers that divide ecoregions, which can overestimate the risk of some short-distance trips within the same ecoregions. Also, the change in shipping patterns due to COVID-19 requires further investigation into changes in shipping behavior and their impact on BWRA, which is a concern for future studies.
@@ Methods
Connected components: A component is a group of vertices that are connected to each other, and a network that has more than one group of vertices that are not connected is called a non-connected graph. In an arbitrary graph, vertices $i$ and $j$ are in the same component $G^{\prime}=\{V^{\prime},E^{\prime}\}$, $V^{\prime}\subseteq V$, if and only if $\{\forall i\in V^{\prime},\forall j\in V^{\prime}|d_{ij}<\infty\}$, meaning that it is possible to travel from any vertex $i$ to any vertex $j$ in a finite number of steps, where $d_{ij}:V\times V\rightarrow\mathbb{R}$ is a function that returns the distance between any two vertices [56]. Connected components are defined for undirected graphs but can also be extended to directed graphs, resulting in weekly and strongly connected components, which help in identifying absorbing regions.
@@
Geodesic distance $(d_{ij}^{E})$: The geodesic distance, as defined in Equation 12, measures the shortest distance between two vertices $i\in V$ and $j\in V$ on a sphere’s surface [57]. This metric incorporates the latitudes $\phi_{i}$ and $\phi_{j}$ of vertices $i$ and $j$, the difference $\bigtriangleup^{\lambda}_{ij}$ between their longitudes $\lambda_{i}$ and $\lambda_{j}$, and the Earth’s radius $\mathbf{R}$ (6,371 km). All values are calculated in radians.
@@
$\hfill d_{ij}^{E}=\left(sin\left(\phi_{i}\right)\times sin\left(\phi_{j}\right)+cos\left(\phi_{i}\right)\times cos\left(\phi_{j}\right)\times cos\left({\bigtriangleup^{\lambda}_{ij}}\right)\times\mathbf{R}\right)\hfill$ (12)
@@
Shortest distance $(d_{ij}^{N})$: A path $S$ between two vertices $i$ and $j$ is a sequence of connected vertices $S^{n}=\left<v_{1},v_{2},\text{...},v_{q-1},v_{q}\right>$, where each consecutive vertex is connected through an edge $S^{n}_{m},S^{n}_{m+1}\right>\in E$ for all $m\in[1,|S^{n}|[$. The shortest directed path $d_{ij}^{N}$ is obtained by minimizing the weight function ${\mathit{f}:E^{n}\rightarrow\mathbb{R}}$, which describes the cost of the paths among all possible paths $\mathbb{S}=\{S^{1},S^{2},\ldots,S^{n}\}$ between vertices $i$ and $j$ [58]. The goal is to find the path with the minimum cost, which is determined by the sum of the weights of the edges. The weight is defined as the straight-line distance between the vertices, such as in Equation 13.
@@
Closeness centrality $(C_{C})$: The Closeness measures how closely a vertex is connected to all the other vertices in a network, as determined by the shortest paths between them (refer to Equation 14). A vertex with higher centrality is considered more central and has a shorter average distance to all other vertices in the network [59]. In a port network, the closeness-central ports are expected to have higher traffic of vessels, providing insight into trading behavior and economic relationships inherent to their country and cities.
@@
Betweenness centrality $(C_{B})$: The Betweenness identifies how often a particular vertex is present in the shortest paths of a graph. A higher metric value indicates that the vertex is present in a larger number of shortest paths. In Equation 16, $\sigma(u,v)$ represents the number of shortest paths between vertices $u$ and $v$, and $\sigma(u,v|i)$ represents the number of those paths that pass through vertex $i$. It is worth noting that there is also the Edge-Betweenness, which assigns a Betweenness value to each edge in a graph. This variant considers $\sigma(u,v|i)$ as the number of shortest paths between vertices $u$ and $v$ that pass through edge $i$.
PageRank $(P_{ij})$: The PageRank is based on a mathematical model known as the stochastic Markovian process. This model defines a probability distribution over a set of states, where the probability of transitioning from one state to another is solely correlated with the state immediately preceding it:
$\hfill P_{ij}=\mathbf{P}(X_{n+1}=j\mid X_{n}=i)\hfill$ (17)
The algorithm assesses the significance of a vertex with respect to the significance of the vertices that are linked to it. This way, it measures the vertex contribution based on the number of outgoing edges each adjacent vertex has, ensuring the uniqueness of the edge. To determine vertex importance, the algorithm calculates the stationary transition probability matrix. The values obtained from this calculation indicate the significance of the vertices based on their access probability. Equation 18 illustrates the stationary matrix, where $\pi^{(n)}$ denotes the probability matrix at time $n$, and $\mathbf{P}$ is the transition probability matrix.
$\hfill\pi^{(n-1)}\mathbf{P}=\pi^{(n)}\hfill$ (18)
Features information: Features incorporated in this study are listed in Table 3, including shipping fluxes, distances, international trade volume, and the graph metrics computed from the global shipping network.
@@
Original port fluxes
Shipping fluxes at source port
Destination region fluxes
Total shipping fluxes at the destination geographical region
Distance
Geodesic distance between the source port to the destination region center
Exportation volume
Exportation volume in US dollars from source to destination
Betweenness centrality
Betweenness centrality at the original port and the median betweenness centrality in the destination region
Closeness centrality
Closeness centrality at the original port and the median closeness centrality in the destination region
Page rank
Page rank at the original port and the median page rank in the destination region

### v3

@@ Discussion
The Transformer Gravity model is built based on a stack of multi-head attention blocks. It incorporates features from the gravity-informed model, including the shipping fluxes, origin and destination locations, and the geographical distance between them. We enhanced the feature set by integrating extra factors, such as bilateral trade information and graph metrics from the global shipping network, to improve the model’s ability to predict ship traffic flows. We validated our approach by comparing it with established gravity models and machine-learning regression techniques to showcase its advancements.
Our findings established the superiority of the Transformer Gravity model across several performance indicators. Notably, the model achieved the highest CPC with minimal variance across different data folds, indicating its robustness and adaptability to new datasets. The best average CPC in the cross-validation set was 0.864 for the 3 and 5-layer models. Besides, the 3-layer one demonstrated the lowest mean error, 0.080, and the highest mean correlation, 0.977. Variations in performance among baseline models indicated their inadequacies in dealing with diverse data scenarios, thus reinforcing our model’s resilience.
The high performance of our Transformer Gravity model for ship traffic flow prediction contributes to a more accurate representation of shipping intensity in the risk assessment of invasive species through BWRA. Despite some discrepancies in the high-risk interval, the evaluation results show an overall low error of $0.208$ and a high similarity between predicted and actual environmental distances among port pairs scaled by ship traffic flows. This has contributed to a more accurate risk assessment for the spread of NIS through ballast water and thus provides a valuable reference for future global ballast water risk management. Specifically, the current Ballast Water Management Convention of the International Maritime Organization (IMO) regulates ship ballast water exchanges, mandating they occur away from coastal areas and at specified depths to decrease the survival and spread of invasive species [40]. To further reduce the survival of NIS, various ballast water treatment methods have been implemented. These include ballast water exchange and on-board ballast water treatment systems (BWTSs) that employ physical, mechanical, and chemical approaches. However, none of these methods are effective at all times [41]. Although IMO regulations became mandatory in 2017, a regional study indicates that as of 2019, less than 20% of ships arriving in the United States and Australia were equipped with BWTSs [42]. This gap in compliance and the partial effectiveness of current ballast water treatments highlight the necessity for advanced predictive models. The Transformer Gravity model plays a pivotal role in finding ports characterized by intense shipping activity through its ability to forecast shipping intensity across diverse maritime pathways. This predictive capability is paramount, especially for routes forecasted to exhibit elevated shipping volumes and a heightened risk of NIS introduction. By identifying these high-risk routes, our model serves as a valuable tool for governmental agencies, urging them to strengthen monitoring and management efforts. This approach is essential to uphold compliance with IMO regulations and mitigate the risks of NIS.
@@ Methods
[…]
Methods
Global Shipping Network Representation
The global shipping network can be described as $G=(V,E,W)$, where $V$ represents the set of ports, $E$ is the set of shipping routes connecting pairs of ports, and $W$ is the collection of edge weights. In this context, each weight in $W$ corresponds to the number of individual trips $T_{ij}$ between ports $i$ and $j$:
$\hfill W=\{w_{ij}:w_{ij}=\sum_{t}T_{ij}(t)\,\forall(i,j)\in E\}\hfill$ (2)
Figure 7: The experiment pipeline for analyzing and predicting links in the global shipping network from 2017 to 2019 includes constructing the global shipping network, calculating centrality and PageRank graph metrics, fully connecting the shipping network for disconnected and weakly connected components, and performing link prediction to identify potential real shipping connections.
Complex Network Metrics Computation
We analyzed the shipping network by calculating various graph metrics for each node. These metrics included betweenness centrality, closeness centrality, and PageRank. The weights of the edges utilized in these metrics computations correspond to the number of trips along these individual shipping routes, as defined in Equation 2. Subsequently, we introduced the graph metrics used for shipping network analysis:
Betweenness centrality $(C_{B})$: The Betweenness identifies how often a particular vertex is present in the shortest paths of a graph. A higher metric value indicates that the vertex is present in a larger number of shortest paths. In Equation 3, $\sigma(u,v)$ represents the number of shortest paths between vertices $u$ and $v$, and $\sigma(u,v|i)$ represents the number of those paths that pass through vertex $i$. The Edge-Betweenness, which assigns a Betweenness value to each edge in a graph, is worth noting. This variant considers $\sigma(u,v|i)$ as the shortest paths between vertices $u$ and $v$ that pass through edge $i$. In the shipping network, betweenness centrality quantifies the frequency with which a port serves as an intermediary on the shortest paths between other ports; ports with high betweenness behave as bridges in the shipping network.
Closeness centrality $(C_{C})$: The Closeness measures how closely a vertex is connected to all the other vertices in a network, as determined by the shortest paths between them (refer to Equation 4). A vertex with higher centrality is considered more central and has a shorter average distance to all other vertices in the network [43]. In a port network, the closeness-central ports are easily accessible from all other ports in the shipping network. They are expected to have higher vessel traffic, providing insight into trading behavior and economic relationships inherent to their country and cities.
PageRank $(P_{ij})$: The PageRank [44] is based on a mathematical model known as the stochastic Markovian process. This model defines a probability distribution over a set of states, where the probability of transitioning from one state to another is solely correlated with the state immediately preceding it:
$\hfill P_{ij}=\mathbf{P}(X_{n+1}=j\mid X_{n}=i)\hfill$ (5)
The algorithm assesses the significance of a vertex concerning the significance of the vertices linked to it. This way, it measures the vertex contribution based on the number of outgoing edges each adjacent vertex has, ensuring the uniqueness of the edge. The algorithm calculates the stationary transition probability matrix to determine vertex importance. The values obtained from this calculation indicate the significance of the vertices based on their access probability. Equation 6 illustrates the stationary matrix, where $\pi^{(n)}$ denotes the probability matrix at time $n$, and $\mathbf{P}$ is the transition probability matrix.
$\hfill\pi^{(n-1)}\mathbf{P}=\pi^{(n)}\hfill$ (6)
As PageRank identifies essential nodes in a graph — ports with high metric values are more influential and likely to be frequently visited by ships from other important ports.
Port Name [36] $C_{B}(i)$ $C_{C}(i)$ $P_{ij}$ Water Body Country Region
As Suways 0.018 0.532 0.003 Gulf of Suez Egypt Northern Africa
Given that betweenness and closeness centrality measures involve the calculation of shortest distances, we also applied the reciprocals of the weights defined in Equation 2 to calculate the graph metrics. This transformation aligns with interpreting weights as distances. However, the centrality rankings of the ports listed in Table 2 remained largely unchanged even after applying the reciprocal weights. This suggests that the topology and connectivity patterns are more critical attributes in the real-world shipping network characterized by sparsity and a highly skewed degree distribution. Furthermore, our shipping network modeling strategy aligns with the principles of gravity models. The strength of relationships is determined by the masses of sources and destinations, with larger objects driving more gravitational pull.
Fully Connected Shipping Network
@@
A component is a group of vertices connected to each other, and a network with more than one group of vertices that are not connected is called a non-connected graph. In an arbitrary graph, vertices $i$ and $j$ are in the same component $G^{\prime}=\{V^{\prime},E^{\prime}\}$, $V^{\prime}\subseteq V$, if and only if $\{\forall i\in V^{\prime},\forall j\in V^{\prime}|d_{ij}<\infty\}$, meaning that it is possible to travel from any vertex $i$ to any vertex $j$ in a finite number of steps, where $d_{ij}:V\times V\rightarrow\mathbb{R}$ is a function that returns the distance between any two vertices [45]. Connected components for directed graphs are divided into weakly and strongly connected components, both of which identify absorbing regions.
When analyzing the shipping network, we observed the presence of 2 disconnected components and 13 components weakly connected (i.e., these subgraphs are unilaterally connected). These disconnected and weakly connected segments pose a challenge for our proposed framework, which relies on integrating features across all possible destinations from each source port. Such disconnections can compromise the model’s ability to generate accurate or well-defined predictions for shipping flows between isolated areas. Thus, we transformed the original network into a complete graph by connecting each pair of nodes $V_{i}$ and $V_{j}$ in the shipping network $G$. We assigned a small weight, $w^{\prime}=0.1$, to these newly established links and denoted the complete graph as $G^{\prime}$. More specifically, we set the weights of these non-existent edges to be inversely proportional to the distance. This approach considers the low probability of vessels leaving one port from a component and directly reaching another, ensuring an accurate representation of the network’s connectivity. This action mitigates the issues arising from data sparsity and establishes a uniform data structure, thereby enhancing the robustness of the flow estimation framework. With the fully connected shipping network $G^{\prime}$, we proceed with the flow estimation framework by forecasting whether a trajectory exists by solving a link prediction problem in $G^{\prime}$, which can be framed as a binary classification task within machine learning models. This action provides concrete source-destination pairs well-prepared for building feature sets and modeling the gravity force.
Network Feature Extraction, Representation, and Purpose
@@
To perform link prediction, we first separated the shipping data from 2019 for testing and retained the data from 2017 and 2018 for training and validation. Then, we prepared features for this binary classification task. We calculated the Haversine distances between every pair of ports, which is defined in Equation 7, measures the shortest distance between two vertices $i\in V$ and $j\in V$ on a sphere’s surface [46]. This metric incorporates the latitudes $\phi_{i}$ and $\phi_{j}$ of vertices $i$ and $j$, the difference $\bigtriangleup^{\phi}_{ij}$ between $\phi_{i}$ and $\phi_{j}$, the difference $\bigtriangleup^{\lambda}_{ij}$ between their longitudes $\lambda_{i}$ and $\lambda_{j}$, and the Earth’s radius $\mathbf{R}$ (6,371 km) — all values are in radians.
@@
$\hfill d_{ij}^{E}=2\times arcsin\left(sin^{2}\left({\bigtriangleup^{\phi}_{ij}}\right)+cos\left(\phi_{i}\right)\times cos\left(\phi_{j}\right)\times sin^{2}\left({\bigtriangleup^{\lambda}_{ij}}\right)\right)\times\mathbf{R}\hfill$ (7)
However, Haversine distances only provide the geodesic approximation and cannot capture the real sea routes that the ships have traveled. Therefore, we also computed the sea-route distances between port pairs. We obtained a more accurate representation using a Python package that models the shortest routes and calculates the sea route distances using historical trajectories [47]. Finally, we used these distances to evaluate the importance level $I_{ij}$ for each edge $\langle i,j\rangle$ in the complete graph $G^{\prime}$. Inspired by straightness centrality measuring the node connectivity by the straightness of the shortest distance [48], this metric combines the normalized flow size $\hat{w}_{ij}$ and the normalized Haversine distance $\hat{d_{ij}^{E}}$, using a small constant $\epsilon$ to prevent division by zero. Connections with high shipping flows and shorter distances are deemed more important:
Following is an explanation of the shortest distance and straightness centrality concepts in network analysis, from which we drew inspiration for our edge importance metric (Equation 8):
@@
Shortest distance $(d_{ij}^{N})$: A path $S$ between two vertices $i$ and $j$ is a sequence of connected vertices $S^{n}=\left<v_{1},v_{2},\text{...},v_{q-1},v_{q}\right>$, where each consecutive vertex is connected through an edge $S^{n}_{m},S^{n}_{m+1}\right>\in E$ for all $m\in[1,|S^{n}|[$. The shortest directed path $d_{ij}^{N}$ is obtained by minimizing the weight function ${\mathit{f}:E^{n}\rightarrow\mathbb{R}}$, which describes the cost of the paths among all possible paths $\mathbb{S}=\{S^{1},S^{2},\ldots,S^{n}\}$ between vertices $i$ and $j$ [49]. The goal is to find the path with the minimum cost, which is determined by the sum of the weights of the edges. The weight is the straight-line distance between the vertices, such as in Equation 9.
@@
Next, we incorporated the Haversine distance (Equation 7), sea route distance [47], and edge importance (Equation 8) into the feature set. We then employed machine learning classification models, including Logistic Regression, KNeighbors, Decision Tree, XGBoost, and Random Forest, to determine the existence of a link between two ports. Upon labeling the real (true class) and pseudo (false class) links in the complete network $G^{\prime}$, we observed a significant class imbalance (2.3% real links and 97.7% pseudo links); consequently, we employed stratified sampling of the pseudo links based on their spatial distribution to balance the number of examples in each class in a manner that preserved the data’s characteristics. The models were engaged in a binary classification task, where we utilized 75% of data from 2017 and 2018 for training and 25% for validation. During the training phase, a 5-fold grid search cross-validation was implemented to fine-tune the hyperparameters of each model. Finally, we tested the models on unseen data from 2019, reinforcing the validity of our approach.
Features used with Gravity-Informed Models
Based on predicted shipping trajectories, we have origin-destination pairs for preparing gravity model features. Features incorporated in gravity-informed models are listed in Table 3, including shipping fluxes, distances, international bilateral trade volume [33], and the centrality graph metrics extracted from the global shipping network. In particular, the bilateral trade volume data is country-based. Therefore, we used the source and destination countries to extract trade volume data and integrate it into the feature set.
@@
Original port fluxes Shipping fluxes at source port
Destination region fluxes Total shipping fluxes at the destination geographical region
Distance Geodesic distance between the source port to the destination region center
Bilateral trading volume Exportation volume in US dollars from source country to destination country
Betweenness centrality Betweenness centrality at the original port and the median betweenness centrality in the destination region
Closeness centrality Closeness centrality at the original port and the median closeness centrality in the destination region
Page rank Page rank at the original port and the median PageRank in the destination region
@@
Transformer Gravity Model
In this study, we incorporate the self-attention mechanism of the Transformer architecture [34] into our proposed framework. Compared to the conventional MLPs structure, the self-attention mechanism can inspect the input sequence and weigh to identify and prioritize the most relevant elements for generating the output. This characteristic enables our model to capture crucial dependencies in vessel mobility effectively flows. Additionally, the self-attention mechanism accomplishes high performance with fewer parameters, making it a computationally efficient model. This section shows how we model the Transformer Gravity, combining the characteristics of the Gravity Model and the self-attention mechanism.
Problem Definition
Based on the pairs of source-destination ports obtained from the previous link prediction step, we encode the destination ports into 17 geographical regions according to the ISO-3166 standard [51], where ports within regions are expected to have a similar set of organisms and, therefore, share similar habitat. Given the limitations in the granularity of our data, which does not extend to cover all geographic areas, this regional categorization makes it feasible to model and incorporate all intra-region locations within the landscape. Over the encoded representation of regions, we now aim to estimate the sizes of shipping mobility flows between each source-destination pair $(P_{i},P_{j})$, where $P_{i}$ is the source port and $P_{j}$ is the destination port that pertains to a unique geographical region. A ship departing from a source port may have one or more destination ports in the same or different regions, and following the Deep Gravity [29] method, the goal is to estimate probabilities of the ships traveling to these geographical regions, becoming a multiclass classification task.
Predict target: We compile a set of 10 features from various sources for each pair $(P_{i},P_{j})$, such as shipping fluxes at ports, geodesic distances between source and destination, bilateral trade volume source from The Atlas of Economic Complexity [33], the graph metrics computed with the global shipping network; detailed information about these extracted features is provided in Table 3 in the Methods. We represent the feature vector for each source-destination pair as $x_{ij}=\left<m_{1},m_{2},\ldots,m_{10}\right>$. Given the ships from each source port travel to multiple destination regions, these feature vectors are aggregated into a single data sample $X_{i}=\{x_{i1},x_{i2},\ldots,x_{iN}\}$, where $N$ is the number of destination regions in the data sample, and $1\leq N\leq 17$. Each destination region is represented as a class, so the prediction has $N$ classes for a sample. Since samples of varying lengths cannot be wrapped to a tensor for batch processing, we set the batch size to 1 for model input. Using $\hat{y}_{ij}$ to represent the estimated size of the mobility flow between $(P_{i},P_{j})$, which is the target of the prediction, we have:
where $O_{i}$ represents the total number of ships departing from source port $i$. $p_{ij}$ is the probability of ships traveling from source port $i$ to the destination region $j$ among all the possible destinations $1\sim N$, and $f(x_{ij})$ is the model output of feature vector $x_{ij}$.
Evaluation metric: The Common Part of Commuters (CPC) [52, 53, 54] is designed to measure the similarity between two sets of data, which could represent various aspects such as the volume of commuters, traffic, or trade between different locations. The metric calculates how much overlap there is between the predicted values $\hat{y}_{ij}$ and the actual values $y_{ij}$. In Equation 13, $M$ represents the number of source ports and $N$ the number of destination regions. The values $\hat{y}_{ij}$ and $y_{ij}$ correspond to the flow of vessels from source port $i$ to destination region $j$, based on a model’s prediction and the actual observed values, respectively. In this case, a high value of the CPC means a significant overlap between the predicted and actual datasets. Specifically, it would indicate that the predictions accurately capture the true data distribution patterns, with most predictive quantities closely matching the actual quantities. Contrarily, a low value of the CPC would suggest that there is little overlap between the predictions and the actual data, indicating that the model’s predictions diverge significantly from the observed data, which could be due to underprediction or overprediction in various parts or a general misalignment of the model with the reality. Accordingly, CPC considers the minimum common value between the predicted and actual data for each pair of source and destination ports, measuring the intersection over the values union:
To evaluate the model’s performance comprehensively, besides the CPC, we included the Normalized Root Mean Square Error ($NRMSE$) — lower is better — and Pearson Correlation Coefficients ($Corr.$) — higher is better — to measure the normalized errors and the correlation between the predictions and observations, and to provide a multi-faceted assessment of the model’s accuracy and reliability, defined as:
Model Framework
Our proposed Transformer Gravity model is composed of three main components: (1) the input embedding layer, which maps the input feature vectors to a higher-dimensional space that is compatible with the Transformer architecture; (2) the multilayer Transformer encoder, which involves the self-attention and feed-forward blocks that process the embeddings to capture complex relationships between input features; and, (3) the output linear layer, which maps the processed embeddings to the target flow predictions, computes loss and CPC and performs backpropagation based on the loss values. The input embedding layer (1) helps us organize and make sense of that data by putting it into a format that is easier for a machine to understand. This format is represented as vectors of numerical values, where each dimension captures specific input data characteristics and is fed into subsequent layers of the model for further processing until achieving the output layer. The output represents the predicted value, which, in our case, is the inflow or outflow of vessels between a pair of ports. The multilayer Transformer encoder (2) looks at all the different parts of information organized in the first component and finds out how they relate. Transformers are known for spotting patterns and connections between different parts of the training data. This helps our model understand the complex relationships between various aspects of the shipping data, like which routes are busiest, or ports are most important. The output linear layer (3) helps us make predictions based on the patterns and connections found by the previous layer. This layer will predict how much shipping traffic will be in a particular area. If the predictions are not entirely correct, this layer also helps us learn from our mistakes and improve our future guesses due to the backpropagation mechanism. Backpropagation is a process where the model adjusts its internal parameters based on the difference between its predictions and the actual outcomes, gradually refining its understanding of the data and improving its predictive power over time. Figure 5 presents the model pipeline using two stacked Transformers modules and provides a glance at the layer’s relationships.
Our experiment defines the number of heads as $h=2$, and the heads run operations in parallel.
Layer Normalization. After the multi-head attention layer, a dropout layer randomly sets a certain percentage of elements to 0. The dropout ratio $p$ is set to $0.1$ as per Equation 18. Next, a skip connection is applied to add the input features $\textbf{z}_{0}$ to the output of the dropout layer $\textbf{z}_{dropout}$ before the self-attention block. This helps retain the input features’ information and prevents vanishing gradients during backpropagation. The output of this connection, $\textbf{z}_{skip}$, is then normalized using Equation 18, where $\mu$ is the mean and $\sigma$ is the standard deviation of $\textbf{z}_{skip}$ with a small bias. The affine parameters $\alpha$ and $\beta$ are initialized as 1 and 0, respectively, and can be optimized during training.
Training and Optimization. Our Transformer Gravity model has three transformer encoder layers stacked together to capture complex input embedding dependencies. Still, the number of stacked layers can be changed to match different requirements and needs. The output value $\textbf{z}_{5}$ is obtained by passing the output of Equation 19 through a second and later third multi-head attention and feed-forward network block. The output value of the model is denoted as $f(x_{ij})$, and it produces a sequence of output values for a single data sample with a length of $N$. This sequence is then applied to a softmax function to produce probabilities $\{p_{i1},p_{i2},\ldots,p_{ij},\ldots,p_{iN}\}$ for $N$ classes. The predicted flow sizes $\{\hat{y}_{i1},\hat{y}_{i2},\ldots,\hat{y}_{ij},\ldots,\hat{y}_{iN}\}$ for each destination are obtained by multiplying these probabilities with the total outflows $O_{i}$ from the source port, as given in Equation 11. The loss for every sample is computed using Equation 12, and these losses are collected to derive the total loss. The model’s parameters are updated with each loss by processing a single sample. The summed CPC across all samples is calculated using Equation 13. After each training epoch, the summed CPC is divided by the number of samples $M$ (i.e., the number of source ports) to obtain the average CPC of that epoch. We used the Adam optimizer with $L_{2}$ regularization on the weights during training. We reduced the learning rate by a factor of 0.1 when there was no improvement in the validation CPC after 10 epochs. We used early stopping when there was no improvement with the validation CPC after 20 epochs to prevent overfitting and improve training efficiency.